Nursing and pharmacy students must solve dosage composition problems similar to this exercise. Set up an equation and solve by an appropriate method.
A doctor prescribes 0.5 liter of a 4% glucose solution. How much distilled water (0% glucose) and 50% glucose solution must be blended to obtain the proper solution?

Answer by richwmiller(17219)

d+g=0.5,
0*d+0.5*g=0.04*0.5
d=0.5-g
0*(0.5-g)+0.5*g=0.02
0.0-0g+0.5*g=0.02
0.5*g=0.02
g=0.04
d=0.5-g
d=0.46 liter of 0% distilled
g=0.04 liter of 50% glucose
check
0*0.46+0.5*0.04=0.04*0.5
0.0+0.02=0.02
0.02=0.02
ok